# Building UD Cairo for Old English in the Classroom

### v1

@@ Abstract
Lauren Levine, Junghyun Min, and Amir Zeldes Linguistics Department, Georgetown University Washington, DC, USA {lel76, jm3743, Amir.Zeldes}@georgetown.edu
@@ 1 Introduction
Treebanking historical languages is a challenging task for multiple reasons – to name just a few: 1. annotators are never native speakers; 2. it is difficult to recruit annotators with sufficient background in the language, in theoretical syntax, and in corpus annotation; 3. guidelines and sample datasets illustrating annotation principles are often tailored to modern languages, with examples that would be implausible in context for ancient texts, and conversely, without examples for contexts that are common in such texts.
@@
In Universal Dependencies (UD, de Marneffe et al. 2021), the most popular treebanking formalism, there is no single template for starting a treebanking project. However, a set of basic sentences is sometimes suggested which is known as the Cairo sentences, originally proposed at CICLing 2015 in Cairo. [^1] https://github.com/UniversalDependencies/cairo The data has been translated and used as a UD starting point for a number of languages, including Ligurian Lusito and Maillard (2021), Luxembourgish Plum et al. (2024) and Gujarati Jobanputra et al. (2024). In this paper, we therefore propose and report on a classroom historical treebanking activity creating a UD Cairo-style corpus for Old English (OE), including generating the text, treebanking proper, and the addition of supplementary annotations. We evaluate treebanking from students with and without previous exposure to UD, and run some parsing experiments to reveal how similar or distinct OE syntax is compared to Modern English (ME). We release all data publicly and hope this work inspires others to incorporate more treebanking into classroom activities.
@@ 2 LLM-assisted UD Cairo in Ald-Ænglisc
Translating the UD Cairo sentences to Old English (Ald-Ænglisc) is a non-trivial task, especially considering the goal of expressing the basic constructions of the language in the most natural possible way, without introducing translation effects conforming to the Modern English syntax or lexicon. As a first step, we considered simply asking an LLM, in this case GPT4o, to translate the sentences (see Appendix A for the full list of sentences with initial GPT outputs and final versions).
@@
The model translated ‘the girl’ into ‘seo mæden’ - a correct lexical choice with an incorrect article gender (feminine, when mæden was in fact neuter in OE). The model also tended to insert indefinite articles (here ‘ǣnne’), which were not yet grammaticalized or expected in OE, where the Modern English original had them, resulting in superfluous ‘ones’ (e.g. ‘one letter’ for ‘a letter’, where just ‘letter’ would be correct). We also note that GPT recreates the input word order with the ‘to’ dative, rather than using the more usual ditransitive construction, (‘wrote her friend a letter’), which was used more in OE. Finally, the word for ‘letter’ is a creative invention, compounded from two attested OE stems, but which is not attested as a word. In the final data, we therefore corrected the gender error, removed the indefinite article, used the ditransitive word order, and used Bosworth Toller’s Anglo-Saxon Online Dictionary [^2] https://bosworthtoller.com/ to select an attested word for ‘letter’, resulting in 2:
@@ 4 Parsing experiments
In order to shed light on how much Old English dependency syntax analysis differs from that of Modern English, we conduct parsing experiments on our OE Cairo sentences using dependency parsers trained on ME. We train DiaParser (Dozat and Manning, 2017), a neural biaffine dependency parser, using English ELECTRA (Clark et al., 2020) and multilingual mBERT (Devlin et al., 2019) embeddings, in order to compare the impact of monolingual vs multilingual pre-trained inputs. For training data, we use version 11 of the GUM corpus Zeldes (2017), which is the largest available UD corpus of Modern English, containing 268k tokens across a variety of genres. We evaluate the ELECTRA and mBERT models on our OE Cairo sentences, using our various token annotations (Form, Lemma, Hyperlemma, and Gloss) as input. The results of these experiments are reported in Table 2, giving the UAS and the LAS for each model. Form refers to the unaltered text of the tokens, and Lemma refers to the uninflected form of the tokens. Hyperlemma and Gloss are as described in Section 3.2. $+$Norm indicates the addition of orthographic normalization to the annotation of the preceding row – replacing OE letters that no longer exist in ME ("th" for "þ", "e" for "æ", etc.).
@@
For both models, performance increases from Form to Hyperlemma and again to Gloss, for both UAS and LAS, with the mBERT model giving the highest score of UAS 85.38% and LAS 78.36%. Using Hyperlemma and Gloss for the model input allows us to abstract away from the lexical differences between Old English and Modern English. Although parsing the Gloss leads to a large improvement in performance over parsing the Form (LAS $\Delta$ 62.65% for the ELECTRA model and LAS $\Delta$ 52.63% for the mBERT model), there is still a large drop in performance when compared to the ME dev data (LAS $\Delta$ 17.73% for ELECTRA and LAS $\Delta$ 13.59% for mBERT). This indicates that there are notable syntactic differences between OE and ME for dependency parsing, even after controlling for lexical differences. Still, the scores for parsing on glossed tokens of Old English are high enough to suggest that leveraging Modern English models to parse glossed Old English texts could be an effective method to bootstrap the creation of Old English syntactic dependency datasets in the future.
@@ Appendix A GPT initial outputs
    “He bought a car but his brother just a bike.”

### v2

@@ Abstract
Lauren Levine, Junghyun Min, and Amir Zeldes Georgetown University, Department of Linguistics {lel76, jm3743, amir.zeldes}@georgetown.edu
@@ 1 Introduction
Treebanking historical languages is a challenging task for multiple reasons – to name just a few: 1. annotators are never native speakers; 2. it is difficult to recruit annotators with sufficient background in the language, in theoretical syntax, and in corpus annotation; 3. guidelines and sample datasets illustrating annotation principles are often tailored to modern languages, with examples that would be implausible in context for ancient texts; 4. there is a scarcity of previous work for annotators to reliably refer to, making examples for contexts that are common in ancient texts difficult to find.
@@
In Universal Dependencies (UD, de Marneffe et al. 2021), the most popular treebanking formalism, there is no single template for starting a treebanking project. However, a set of basic sentences is sometimes suggested, known as the Cairo sentences and originally proposed at CICLing 2015 in Cairo. [^1] https://github.com/UniversalDependencies/cairo The data has been translated and used as a UD starting point for a number of languages, including Ligurian Lusito and Maillard (2021), Luxembourgish Plum et al. (2024) and Gujarati Jobanputra et al. (2024). In this paper, we therefore propose and report on a classroom historical treebanking activity creating a UD Cairo-style corpus for Old English (Ald-Ænglisc or OE), including generating the text, treebanking proper, and adding supplementary annotations.
Although some work on treebanking OE and later stages of older English exists – notably The York–Toronto–Helsinki Parsed Corpus of Old English Prose (YCOE; Taylor, 2007), the Penn-Helsinki Parsed Corpus of Early Modern English (PPCEME; Kroch et al., 2004), and their follow-up work (Kroch, 2020; Kulick et al., 2023) – no such data exists in UD. We add to the literature by providing the first OE contribution to UD. We evaluate treebanking from students with and without previous exposure to UD, and run some parsing experiments to reveal how similar or distinct OE syntax is compared to Modern English (ME). We release all data publicly [^2] https://universaldependencies.org/treebanks/ang_cairo/index.html and hope this work inspires others to incorporate more treebanking into classroom activities.
@@ 2 LLM-assisted UD Cairo in Ald-Ænglisc
Translating the UD Cairo sentences to Old English is a non-trivial task, especially considering the goal of expressing the basic constructions of the language in the most natural possible way, without introducing translation effects conforming to the Modern English syntax or lexicon. As a first step, we considered simply asking an LLM, in this case GPT4o, to translate the sentences (see Appendix A for the full list of sentences with initial GPT outputs and final versions).
@@
The model translated ‘the girl’ into ‘seo mæden’ - a correct lexical choice with an incorrect article gender (feminine, when ‘mæden’ was in fact neuter in OE). The model also tended to insert indefinite articles (here ‘ǣnne’), which were not yet grammaticalized or expected in OE, where the Modern English original had them, resulting in superfluous ‘ones’ (e.g. ‘one letter’ for ‘a letter’, where just ‘letter’ would be correct). We also note that GPT recreates the input word order with the ‘to’ dative, rather than using the more usual ditransitive construction, (‘wrote her friend a letter’), which was used more in OE. Finally, the word for ‘letter’ is a creative invention, compounded from two attested OE stems, but which is not attested as a word. In the final data, we therefore corrected the gender error, removed the indefinite article, used the ditransitive word order, and used Bosworth Toller’s Anglo-Saxon Online Dictionary [^3] https://bosworthtoller.com/ to select an attested word for ‘letter’, resulting in 2:
@@ 3.1 UD trees and classroom annotation
3.2 Disagreement Analysis
Figure 2: Confusion matrix for gold dependency relation (deprel) labels vs the student annotations. Deprel subtypes are collapsed, only deprel labels that occur > 5 times are included, and punct was excluded.
In Figure 2, we show a confusion matrix comparing our final gold labels for the Cairo UD sentences with the student annotations from our classroom activity. In line with our high inter-annotator agreement scores in the previous section, we see that there is strong overlap in the selection of dependency relation labels. However, there is still a smattering of disagreements between the gold labels and the student labels. The dependency relation that was mislabeled by students the most was nmod, which was mislabeled 6 times (3 times as det, 2 times as obj, and 1 time as nsubj). These confusions primarily stemmed from students not understanding when to apply the nmod:poss subtype when possessive case was involved in the sentence. The label that was most misapplied by students was obj, which was misapplied a total of 9 times (2 times each on nmod, obl, and orphan relations, and 1 time each as on conj, det, and xcomp relations). In Figure 3, we show a dependency analysis for one of the UD Cairo sentences where a student both misapplied the obj label rather than utilizing the orphan relation, as well as incorrectly using det for a possessive pronoun. We see that the student uses a promotion strategy, rather than utilizing orphan. Likely the use of more complicated relations, such as orphan, could be learned sufficiently with more time devoted to annotator training.
{dependency}
[arc edge, arc angle=80, text only label, label style=above] {deptext}[column sep=.7cm] He & bohte & cræt & ac & his & broþor & hweol-bearwe
He & bought & car & but & his & brother & wheelbarrow
[…]
Figure 3: Dependency analysis of a ellipsis construction in OE: He bought a car but his brother just a wheelbarrow. ME gloss is included below the OE tokens. The gold analysis is shown on top in black, and a student analysis is shown on the bottom in blue.
@@ 4 Parsing experiments
In order to shed light on how much Old English dependency syntax analysis differs from that of Modern English, we conduct parsing experiments on our OE Cairo sentences using dependency parsers trained on ME. While it may be possible and potentially effective to use dependency parsers trained on other modern and historical Germanic languages, we limit the scope our parsing experiments to English. We train DiaParser (Dozat and Manning, 2017), a neural biaffine dependency parser, using English ELECTRA (Clark et al., 2020) and multilingual mBERT (Devlin et al., 2019) embeddings, in order to compare the impact of monolingual vs multilingual pre-trained inputs. For training data, we use version 11 of the GUM corpus Zeldes (2017), which is the largest available UD corpus of Modern English, containing 268k tokens across a variety of genres. We evaluate the ELECTRA and mBERT models on our OE Cairo sentences, using our various token annotations (Form, Lemma, Hyperlemma, and Gloss) as input. The results of these experiments are reported in Table 2, giving the UAS and the LAS for each model. Form refers to the unaltered text of the tokens, and Lemma refers to the uninflected form of the tokens. Hyperlemma and Gloss are as described in Section 3.3. $+$Norm indicates the addition of orthographic normalization to the annotation of the preceding row – replacing OE letters that no longer exist in ME ("th" for "þ", "e" for "æ", etc.).
@@
For both models, performance increases from Form to Hyperlemma and again to Gloss, for both UAS and LAS, with the mBERT model giving the highest score of UAS 85.38% and LAS 78.36%. Using Hyperlemma and Gloss for the model input allows us to abstract away from the lexical differences between Old English and Modern English. Although parsing the Gloss leads to a large improvement in performance over parsing the Form (LAS $\Delta$ 62.65% for the ELECTRA model and LAS $\Delta$ 52.63% for the mBERT model), there is still a large drop in performance when compared to the ME dev data (LAS $\Delta$ 17.73% for ELECTRA and LAS $\Delta$ 13.59% for mBERT). Although some degradation can be expected in cross-corpus parsing Zeldes and Schneider (2023), its magnitude indicates that there are notable syntactic differences between OE and ME for dependency parsing, even after controlling for lexical differences. Still, the scores for parsing on glossed tokens of Old English are high enough to suggest that leveraging Modern English models to parse glossed Old English texts could be an effective method to bootstrap the creation of Old English syntactic dependency datasets in the future.
@@ Appendix A GPT initial outputs
    “He bought a car but his brother just a wheelbarrow.”